OpenStudy (anonymous):

Find the slope of every line that is parallel to the graph of the equation. x+8y=5

OpenStudy (anonymous):

parallel means they have the same slope. the slope of \[ax+by=c\] is \[m=\frac{-a}{b}\] in this case the slope is \[-\frac{1}{8}\]

OpenStudy (angela210793):

8y=5-x y=5/8-x/8 so slope = -1/8

OpenStudy (anonymous):

ohhh i get it thanks